# Diffusion as Shader: 3D-aware Video Diffusion for Versatile Video Generation Control

### v1

@@ 2.2. Controlled video generation
Object manipulation. Object manipulation refers to versatile object movement control for image-to-video generation. Different from camera control, which focuses on changes in perspective, object manipulation emphasizes the movement of the objects themselves. Currently, mainstream methods (Chen et al., 2023a; Li et al., 2024; Mou et al., 2024; Teng et al., 2023; Wang et al., 2024c; Yin et al., 2023; Jain et al., 2024; Ma et al., 2024b; Qiu et al., 2024; Wang et al., 2024f; Yang et al., 2024a; Geng et al., 2024) typically achieve object manipulation by utilizing directed trajectories or modeling the relationships between bounding boxes with specific semantic meanings. However, these methods primarily rely on 2D guidance to represent the spatial movement of target objects, which often fails to accurately capture user intent and frequently results in distorted outputs. ObjCtrl-2.5D (Wang et al., 2024a) tries to address this limitation by extending 2D trajectories with depth information, creating a single 3D trajectory as the control signal. Better than the single 3D trajectory, our method leverages 3D tracking videos, which offer greater details and more effectively represent the motion relationships between foreground and background. This approach enables more precise and realistic object manipulation.
Concurrent works. Recently, several works (Geng et al., 2024; Niu et al., 2024; Koroglu et al., 2024; Jeong et al., 2024; Shi et al., 2024; Lei et al., 2024; Feng et al., 2024a) have explored utilizing motion as control signals. These approaches can be broadly categorized into two groups: 2D motion-based and 3D motion-based methods. (Koroglu et al., 2024; Shi et al., 2024; Lei et al., 2024) leverage 2D optical flow to condition motion, while (Niu et al., 2024; Geng et al., 2024; Jeong et al., 2024) utilize 2D tracks, which are sparser than optical flow, to track or control video motion. In contrast to these methods that rely on 2D motion as guidance,  (Feng et al., 2024a) lifts videos into 3D space and extracts the motion of 3D points, enabling a more accurate capture of spatial relationships between objects and supporting tasks such as object manipulation and camera control. Our method, DaS, also leverages recent tracking methods (Xiao et al., 2024b; Zhang et al., 2025) to construct videos. However, we extend the applicability by unifying a broader range of control tasks, including mesh-to-video generation and motion transfer.

### v2

@@ 2.2. Controlled video generation
Object manipulation. Object manipulation refers to versatile object movement control for image-to-video generation. Different from camera control, which focuses on changes in perspective, object manipulation emphasizes the movement of the objects themselves. Currently, mainstream methods (Chen et al., 2023a; Li et al., 2024; Mou et al., 2024; Teng et al., 2023; Wang et al., 2024c; Yin et al., 2023; Jain et al., 2024; Ma et al., 2024b; Qiu et al., 2024; Wang et al., 2024f; Yang et al., 2024a; Geng et al., 2024) typically achieve object manipulation by utilizing directed trajectories or modeling the relationships between bounding boxes with specific semantic meanings. However, these methods primarily rely on 2D guidance to represent the spatial movement of target objects, which often fails to accurately capture user intent and frequently results in distorted outputs. ObjCtrl-2.5D (Wang et al., 2024a) tries to address this limitation by extending 2D trajectories with depth information, creating a single 3D trajectory as the control signal. Better than the single 3D trajectory, our method leverages 3D tracking videos, which offer greater details and more effectively represent the motion relationships between foreground and background for more precise and realistic object manipulation.
Concurrent works. Recently, several works (Geng et al., 2024; Niu et al., 2024; Koroglu et al., 2024; Jeong et al., 2024; Shi et al., 2024; Lei et al., 2024; Feng et al., 2024a; Zhang et al., 2024) have explored utilizing motion as control signals. These approaches can be broadly categorized into two groups: 2D motion-based and 3D motion-based methods. (Koroglu et al., 2024; Shi et al., 2024; Lei et al., 2024) leverage 2D optical flow to condition motion, while (Niu et al., 2024; Geng et al., 2024; Jeong et al., 2024) utilize 2D tracks, which are sparser than optical flow, to track or control video motion. (Zhang et al., 2024) learns to generate 3D coordinates in the video diffusion model, which 3D awareness. (Feng et al., 2024a) lifts videos into 3D space and extracts the motion of 3D points, enabling a more accurate capture of spatial relationships between objects and supporting tasks such as object manipulation and camera control. Our method, DaS, also leverages recent tracking methods (Xiao et al., 2024b; Zhang et al., 2025) to construct videos. However, we extend the applicability by unifying a broader range of control tasks, including mesh-to-video generation and motion transfer.